OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point shown and is parallel to the line shown below.

OpenStudy (anonymous):

y = 3x – 14 y = 1/3x + 2 y = –3x + 22 y = –1/3x + 8

OpenStudy (anonymous):

It's y = 1/3x + 2, right?

OpenStudy (hba):

Take any two points on the slope,find rise run and then put it in the equation and find it.

OpenStudy (anonymous):

So if I take (0,-3) and (3,2), I just add them into X and Y of the equations?

OpenStudy (hba):

First find slope and then put it in the equation

OpenStudy (anonymous):

The slope is 1/3?? Right?

OpenStudy (hba):

No m=y2-y1/x2-x1

OpenStudy (anonymous):

I got 5/3??

OpenStudy (hba):

Yeah right :) Now put it in the equation :)

OpenStudy (anonymous):

Would I out that under the point slope formula as in y-y1=m(x-x1) and then put that to a slope intercept form?

OpenStudy (anonymous):

judging from the line, the slope of the given graph is 1/3, not 5/3 if the new line passes through (6, 4), then use the point-slope formula with that point and the slope 1/3 y - 4 = (1/3) ( x - 6 ) y = (1/3) x - 2 + 4 y = (1/3) x + 2

OpenStudy (anonymous):

thank you so much!!